Word Problem
Janice has enough money to purchase pebbles to cover 74 square feet. She wishes to form a border aroud her pool. The pool measures 25ft in length and 10 ft in width. How wide should she make the border?

Answer by edjones(8007)

P=2(25+10)=70ft
let x = width length=perimeter
A=L*W= 70x=74 sq ft
x=74/70=1.057 ft wide.
Ed